# The AI Community Building the Future? A Quantitative Analysis of Development Activity on Hugging Face Hub

### v1

@@ Abstract
Open source developers have emerged as key actors in the political economy of artificial intelligence (AI), with open model development being recognised as an alternative to closed-source AI development. However, we still have a limited understanding of collaborative practices in open source AI. This paper responds to this gap with a three-part quantitative analysis of development activity on the Hugging Face (HF) Hub, a popular platform for building, sharing, and demonstrating models. First, we find that various types of activity across 348,181 model, 65,761 dataset, and 156,642 space repositories exhibit right-skewed distributions. Activity is extremely imbalanced between repositories; for example, over 70% of models have 0 downloads, while 1% account for 99% of downloads. Second, we analyse a snapshot of the social network structure of collaboration on models, finding that the community has a core-periphery structure, with a core of prolific developers and a majority of isolate developers (89%). Upon removing isolates, collaboration is characterised by high reciprocity regardless of developers’ network positions. Third, we examine model adoption through the lens of model usage in spaces, finding that a minority of models, developed by a handful of companies, are widely used on the HF Hub. Overall, we find that various types of activity on the HF Hub are characterised by Pareto distributions, congruent with prior observations about OSS development patterns on platforms like GitHub. We conclude with a discussion of the implications of the findings and recommendations for (open source) AI researchers, developers, and policymakers.
@@ 1 Introduction
Open source developers have become central actors in the political economy of artificial intelligence (AI). The rise of open source AI, specifically the practice of releasing and fine-tuning pre-trained models that are freely available [^1] N.B. We use “open models” because they do not meet the requirements to be considered “open source” [1, 2]., has extended open science practices crucial to AI advances [3, 4, 5], including open source software (OSS) [^2] As defined by the OSI [2], OSS is software source code that anyone can inspect, use, modify, or redistribute. and open access to research papers [6] and datasets [7, 8, 9]. Open source AI has attracted attention as a potential challenger to the dominance of startups and Big Tech in AI research and development (R&D) [10, 11]. Grassroots initiatives like EleutherAI [12], BigScience [13], and BigCode [14] have shown the feasibility of open model development [15], while the Hugging Face (HF) Hub has emerged as a popular platform used by millions to host, download, and collaborate on a growing number of models, datasets, and spaces (i.e., web applications to demonstrate and try out models) [16].
While the benefits and risks of open source AI have been widely debated [17, 18, 19, 20], the practices and processes involved in open model development have received relatively little attention. To date, only a handful of scholars have explored various aspects of open model development, including user contributions to grassroots initiatives [13, 15], the role of commercial participation [15, 21], and model maintenance [22]. We contribute to this nascent research agenda with a three-part quantitative analysis of development activity on the HF Hub. First, we investigate typical patterns of various types of activity on the HF Hub in 348,181 model, 65,761 dataset, and 156,642 space repositories (RQ1). Subsequently, we apply social network analysis (SNA) of code contributions to model repositories to investigate the social network structure of the developer community as well as collaboration practices amongst developers (RQ2). We replicate this analysis for models in the sub-fields of natural language processing (NLP), computer vision (CV), and multimodal (MM) for comparative analysis. Finally, we quantify model adoption through the lens of model usage in spaces on the HF Hub (RQ3), providing insights into the widespread use of a minority of models in the HF Hub developer community as well as the key actors driving the development of the most widely used open models.
@@ 2 Related Work
The proliferation of open models, especially foundation models, has ignited heated debate about their potential benefits and risks [19, 17, 18, 42, 20]. On the one hand, open models are said to promise benefits for research, innovation, and competition by lowering entry barriers and widening access to state-of-the-art AI [43]. Drawing on Linus’ Law from OSS development that “given enough eyeballs, all bugs are shallow” [44], proponents argue that open model development and auditing offers safety advantages [45]. In addition, open access to models lowers the barriers for adaptability and customisation for diverse language contexts [46, 19]. On the other hand, open models can pose risks of harm by both well-intended and malicious actors, including the creation of deepfakes [47, 48, 49], disinformation [50, 51], and malware [52, 53]. A study by 25 experts concluded that open models have five distinctive properties that present both benefits and risks: broader access, greater customisability, local adaptation and inference ability, the inability to rescind model access, and the inability to monitor or moderate model usage [19].
@@
It is important to note that a myriad of meanings are attached to “open models” and “open source AI”. Oftentimes it is understood as making pre-trained models, parameters (or “weights”), and documentation available on platforms like the HF Hub. In some cases, it refers to the open collaboration employed to build open models [15]. The description of open models as “open source” has been fiercely contested for failing to meet OSS standards as defined by the OSI [2, 4, 61, 62]. For example, a review found that, “[W]hile there is a fast-growing list of projects billing themselves as ’open source’, many inherit undocumented data of dubious legality, few share the all-important instruction-tuning (a key site where human annotation labour is involved), and careful scientific documentation is exceedingly rare” [63]. Furthermore, companies have been criticised for “open-washing” by describing their models as "open source," to disguise their strategies to set open standards and benefit from crowdsourced contributions [21, 60, 64, 11]. For example, when Meta imposed limits on use of LLaMA 2, Stefano Maffulli from the OSI commented, “Unfortunately, the tech giant has created the misunderstanding that LLaMA 2 is ‘open source’ – it is not. Meta is confusing ’open source’ with ‘resources available to some users under some conditions,’ [which are] two very different things” [61].
@@ 2.3 Learning from Prior Work on OSS development
The various social structures of OSS communities are shaped, amongst others, by the diverse incentives of individuals and companies that participate in OSS development [87, 88]. Individual developers are typically motivated by factors such as personal values, altruism, enjoyment, reputation-building, and career benefits [89, 90, 91, 92]. However, there are also major barriers to participation, including gender disparities [93, 94] and geographic inequalities [84, 86]. Activity tends to be concentrated in the Global North [95] and the English lingua franca is a barrier for many developers [86, 96]. Furthermore, the incentives of OSS developers vary by geography: while developers in the USA show a relatively strong interest in “geek culture”, developers in India and China tend to be motivated primarily by career benefits [97]. Thus, “researchers studying open source should be mindful of geographic variation in what motivates participation and what forms participation may take, particularly outside of the code repository” [84].
Meanwhile companies primarily participate in OSS development for economic and strategic reasons, such as recruiting developers [98, 64, 99], reducing costs [100, 101, 64], influencing OSS projects [102, 101], promoting open standards [100, 103, 104], and building a reputation as an OSS patron [88, 105, 31]. Commercial participation has mixed effects on the social structure of OSS communities. Typically, one company or a few companies emerge as dominant contributors in projects [27, 106]. The dominance of a company is negatively associated with the participation of volunteers, while it is positively associated with the productivity of contributors and the quality of issue reports [107, 72]. It is also common for companies, which may be market rivals, to collaborate in OSS ecosystems [108, 109, 106, 110, 111], which has turned many OSS communities “from networks of individuals into networks of companies” [98].
@@ 3.5 Data Analysis
To investigate development activity on the HF Hub (RQ1), we conducted a descriptive analysis of various types of activity in 348,181 model repositories, 65,761 dataset repositories, and 156,642 space repositories. Pearson correlation coefficients were calculated to assess the pairwise relationships between the activity variables. In addition, we employed the Mann-Whitney U test to compare the levels of activity across repositories with different licences (Permissive, Restrictive, and No licence). The Mann-Whitney U test is a non-parametric test that examines whether two independent samples come from the same distribution, which does not require the data to be normally distributed or to meet the assumption of homogeneity of variance [124]. Given the large sample sizes, the U values are expected to be large, and the salient test statistic is the p-value which indicates the statistical significance of observed differences. Due to capacity constraints in labelling licences, we limited this analysis to repositories with licences used in at least 100 repositories ($n=$339,502, 98% of all repositories). Subsequently, we analysed a snapshot of the social network structure of collaboration on the HF Hub (RQ2), using techniques defined in Table 1 in Appendix A. This analysis provides insights into collaboration patterns in model repositories at this point in time. Furthermore, we analysed collaboration patterns in the three AI sub-fields (NLP, CV, and MM) to enable comparisons. Lastly, we examined model adoption on the HF Hub (RQ3) by calculating the ranked degree of models in the bipartite model usage networks and ranked degree of models in the model co-usage networks to identify the most used models in spaces and their respective developers. These two complementary approaches quantified model popularity (i.e. which models are most frequently used in spaces) and model co-popularity (i.e. which models are most commonly used in conjunction with other models). We replicated this analysis for spaces with NLP, CV, and MM tags for comparative analysis of the three AI sub-fields.
@@ 4.1 Development Activity on the HF Hub
In this section, we present the findings of development activity in the repositories of 348,181 models, 65,761 datasets, and 156,642 spaces on the HF Hub. We present three key findings: right-skewed distributions across different types of activity (Section 4.1.1), strong correlations between development activities (Section 4.1.2), and a significant lack of licences in model and dataset repositories (Section 4.1.3).
@@ 4.1.1 Right-Skewed Distributions in development Activity
Activity per repository is extremely imbalanced. We observe right-skewed distributions of n_likes, n_discussions, n_commits, and n_downloads across the model, dataset, and space repositories (see Figure 1). For example, while the maximum number of likes amongst models is over 9,000, the average model only receives 1.14 likes (see Table 4). The majority of repositories receive minimal engagement. For example, 91% of models and 88% of datasets have 0 likes; 84% of models, 91% of datasets, and 96% of spaces have 0 discussions; and 71% of models and 70% of datasets have 0 downloads. Meanwhile, most activity is concentrated in a small number of repositories. For example, $<1$% of models account for 80% of likes, 10% for 80% discussions, 30% for 80% commits, and $<1$% for 80% downloads. Upon increasing the threshold, 8% of models account for 99% likes, 15% for 99% discussions, and 1% for 99% downloads.
@@
Figure 1: Distributions of Development Activity in Repositories on HF Hub
@@ 4.1.2 Correlation between Community Size and Engagement
4.1.3 Impact of Licences on Collaboration
A significant proportion of model and dataset repositories lack licences, which can create uncertainty and potential legal issues for users and developers. Specifying a licence is not the norm: the majority of model repositories (65%) and datasets (72%) do not have a licence. Amongst the licenced models, the most commonly used licences are Apache v2.0 (37%), MIT (17%), OpenRAIL (14%), and CreativeML OpenRAIL-M (10%). The most used licences for datasets are MIT (28%), Apache v2.0 (15%), OpenRAIL (9%), and licences from the family of Creative Commons v4.0 (7%).
The choice of licence matters: there is a moderate to strong correlation between the use of a licence and level of activity in model repositories (see Figure 3). Furthermore, the Mann-Whitney U tests provide strong evidence of statistically significant differences between collaboration dynamics in model repositories with different types of licences (all tests have $p<0.001$). Specifically, model repositories with permissive licences consistently have the highest levels of activity compared to model repositories with no licence and those with restrictive licences (see Table 5). However, repositories with restrictive licences also exhibit significantly higher activity than those with no licence. This pattern holds across all activity metrics measured, suggesting that while permissive licences foster the highest engagement, restrictive licences also promote more collaboration compared to model repositories that do not have a licence.
@@ 4.1.3 Impact of Licences on Collaboration
Figure 3: Correlations of Activity in Model Repositories with Different Licences
@@ 4.2.1 Collaboration in Model Repositories on the HF Hub
Collaboration is characterised by high reciprocity values, ranging from 0.81 to 1.00 across all k-core levels (see Table 9), indicating the prevalence of mutual relationships amongst developers. The low assortativity values, ranging from -0.49 to 0.08, suggest that developers collaborate regardless of their centrality in the network, implying that other factors, such as shared interests, skills, or project roles, may be more significant in driving collaboration than their network centrality. Furthermore, the relatively low average rich club coefficients, ranging from 0.04 to 0.41, indicate that highly central developers do not primarily collaborate with each other, further emphasising that collaboration is not solely driven by the network centrality of developers.
@@ 5.1.1 Contributions to Academic Literature
Impact of licence on collaboration: The Mann-Whitney U tests show that licence choice significantly impacts the level of activity and engagement in repositories, with permissive licences exhibiting the highest activity levels, followed by repositories with restrictive licences, and finally ones with no licence. Furthermore, the Pearson correlations indicate that the use of a licence (permissive or restrictive) is associated with stronger correlations between various types activity compared to repositories without licences. These findings highlight the important role of licencing decisions in influencing the collaborative and community dynamics in open model development and open source AI projects.
@@ 5.1.3 Recommendations for Future Research
1. Implications of concentrations in the HF Hub developer community: Our analysis confirmed prior observations that the models of a handful of industry-leading companies are dominant amongst the HF Hub developer community [21]. We encourage future research to investigate what these concentrations mean in practice. This includes examining the potential benefits that developers of these popular models may accrue from the model ecosystems, such as increased visibility, crowdsourced contributions (e.g., via commits and discussions), and access to diverse fine-tuned versions shared by other developers on the HF Hub. Furthermore, there is a concern that dominant companies may benefit from developers being locked-in to their ecosystems, potentially limiting competition and entrenching already dominant actors. Future research should investigate the factors contributing to this concentration, such as the reputation of the organisations developing the models, their access to resources and support, or the perceived performance and versatility of their models. Additionally, it would be valuable to examine the broader implications of these concentrations for the AI community, including the impact on research, innovation, and the distribution of benefits and resources.
2. Collaboration in repositories with large communities: Extending our analysis of collaboration across all model repositories on the HF Hub, future research could investigate the social network structure of collaboration in the repositories of models and datasets that have large communities, providing insights into the social practices and processes involved in such collaborations. Furthermore, future research may also compare collaboration practices with regards discussions and commits, building on our observation of differences in the count of unique developers that participate in discussions versus those who make commits (e.g., model repositories have up to 240 discussion starters but only up to 18 committers). Future analyses should also take into account temporal dynamics in this fast-moving field. For example, a temporal analysis of collaboration in the open source AI ecosystem can provide insights into the evolving social structure, trends, and dynamics of open model ecosystems and their communities. In turn, these insights can inform strategies aimed at fostering sustainable growth, innovation, and governance in said ecosystem.
@@
3. Incentives and modes of participation: Future research could investigate the incentives of individual developers and companies. A number of companies have released open models on the HF Hub, such as Meta’s LlaMA models [125], Mistral AI’s Mixtral models [126], and OpenAI’s Whisper models [127]. Often these releases are presented as acts of “AI democratisation” [128]. Future research could critically examine the commercial incentives behind these releases. In addition, future research could examine commercial approaches to model governance and maintenance—for example, if and how companies welcome or engage with community contributions—and if and how companies collaborate with each other on open model development, as they do in OSS development [108, 109, 106, 110, 111, 98].
@@
5. Collaboration beyond the HF Hub: While this analysis provides insights into the community that shares and fine-tunes open models on the HF Hub, we still have a limited understanding of the development of open models that largely takes place in proprietary settings or in the case of collaborative initiatives on other platforms, such as GitHub [15]. Therefore, future research should examine how the HF Hub is used in the wider ecosystem of online platforms as well as offline venues for the collaborative development of open models and datasets. This research direction would enable comparisons of the collaboration patterns amongst model developers and model fine-tuners. Future research could also undertake a multi-sited analysis, examining collaboration on the same project across platforms.

### v2

@@ Abstract
Open model developers have emerged as key actors in the political economy of artificial intelligence (AI), but we still have a limited understanding of collaborative practices in the open AI ecosystem. This paper responds to this gap with a three-part quantitative analysis of development activity on the Hugging Face (HF) Hub, a popular platform for building, sharing, and demonstrating models. First, various types of activity across 348,181 model, 65,761 dataset, and 156,642 space repositories exhibit right-skewed distributions. Activity is extremely imbalanced between repositories; for example, over 70% of models have 0 downloads, while 1% account for 99% of downloads. Furthermore, licenses matter: there are statistically significant differences in collaboration patterns in model repositories with permissive, restrictive, and no licenses. Second, we analyse a snapshot of the social network structure of collaboration in model repositories, finding that the community has a core-periphery structure, with a core of prolific developers and a majority of isolate developers (89%). Upon removing the isolates from the network, collaboration is characterised by high reciprocity regardless of developers’ network positions. Third, we examine model adoption through the lens of model usage in spaces, finding that a minority of models, developed by a handful of companies, are widely used on the HF Hub. Overall, activity on the HF Hub is characterised by Pareto distributions, congruent with OSS development patterns on platforms like GitHub. We conclude with recommendations for researchers, companies, and policymakers to advance our understanding of open AI development.
@@ 1 Introduction
Open source developers have become central actors in the political economy of artificial intelligence (AI). The rise of open source AI, specifically the practice of releasing and fine-tuning pre-trained models that are freely available [^1] N.B. We use “open models” because they do not meet the requirements to be considered “open source” [1, 2]., has extended open science practices crucial to AI advances [3, 4], including open source software (OSS) [^2] As defined by the OSI [2], OSS is software source code that anyone can inspect, use, modify, or redistribute. and open access to research papers [5] and datasets [6, 7, 8]. Open source AI has attracted attention as a potential challenger to the dominance of startups and Big Tech in AI research and development (R&D) [9, 10]. Grassroots initiatives like EleutherAI [11], BigScience [12], and BigCode [13] have shown the feasibility of open model development [14], while the Hugging Face (HF) Hub has emerged as a popular platform used by millions to host, download, and collaborate on a growing number of models, datasets, and spaces (i.e., web applications to demonstrate and try out models) [15].
While the benefits and risks of open source AI have been widely debated [16, 17, 18, 19, 20], the practices and processes involved in open model development have received relatively little attention. To date, only a handful of scholars have explored various aspects of open model development, including user contributions to grassroots initiatives [12, 14], commercial participation in model development [14, 21], model maintenance [22], and the processes and tools used by adjacent open data engineering communities [23].
We contribute to this nascent research agenda with a three-part quantitative analysis of development activity on the HF Hub. First, we investigate typical patterns of various types of activity on the HF Hub in 348,181 model, 65,761 dataset, and 156,642 space repositories (RQ1). Subsequently, we apply social network analysis (SNA) of code contributions to model repositories to investigate the social network structure of the developer community as well as collaboration practices amongst developers (RQ2). We replicate this analysis for models in the sub-fields of natural language processing (NLP), computer vision (CV), and multimodal (MM) for comparative analysis. Finally, we quantify model adoption through the lens of model usage in spaces on the HF Hub (RQ3), providing insights into the widespread use of a minority of models in the HF Hub developer community and the key actors driving their development.
@@ 2 Related Work
The proliferation of open models, especially foundation models, has ignited heated debate about their potential benefits and risks [18, 16, 17, 43, 19, 20]. On the one hand, open models are said to promise benefits for research, innovation, and competition by lowering entry barriers and widening access to state-of-the-art AI [44]. Drawing on Linus’ Law from OSS development that “given enough eyeballs, all bugs are shallow” [45], proponents argue that open model development and auditing offers safety advantages [46]. In addition, open access to models lowers the barriers for adaptability and customisation for diverse language contexts [47, 18]. On the other hand, open models can pose risks of harm by both well-intended and malicious actors, including the creation of deepfakes [48, 49, 50], disinformation [51, 52], and malware [53, 54]. A study by 25 experts concluded that open models have five distinctive properties that present both benefits and risks: broader access, greater customisability, local adaptation and inference ability, the inability to rescind model access, and the inability to monitor or moderate model usage [18].
@@
A myriad of meanings are attached to “open models” and “open source AI”. Oftentimes these terms are understood as making pre-trained models, parameters (or “weights”), and documentation available on platforms like the HF Hub. In some cases, they refer to open collaboration on the development of models [14]. The description of open models as “open source” has been fiercely contested for failing to meet OSS standards as defined by the OSI [2, 4, 63, 64]. For example, when Meta imposed limits on use of LLaMA 2, Stefano Maffulli from the OSI commented, “Unfortunately, the tech giant has created the misunderstanding that LLaMA 2 is ‘open source’ – it is not. Meta is confusing ’open source’ with ‘resources available to some users under some conditions,’ [which are] two very different things” [63].
Companies have been criticised for “open-washing” by promoting their models as “open source” models, when they are typically “open weight” models at most, as a commercial strategy to present themselves as patrons of the digital commons, whilst disguising their intent to set open standards and benefit from crowdsourced innovation [21, 62, 65, 66]. A review of the openness of LLMs found that, “[W]hile there is a fast-growing list of projects billing themselves as ’open source’, many inherit undocumented data of dubious legality, few share the all-important instruction-tuning (a key site where human annotation labour is involved), and careful scientific documentation is exceedingly rare” [66].
@@ 2.3 Learning from Prior Work on OSS development
The various social structures of OSS communities are shaped, amongst others, by the diverse incentives of individuals and companies that participate in OSS development [89, 90, 91]. Individual developers are typically motivated by factors such as personal values, altruism, enjoyment, reputation-building, and career benefits [92, 93, 94, 95]. However, there are also major barriers to participation, including gender disparities [96, 97] and geographic inequalities [86, 88]. Activity tends to be concentrated in the Global North [98] and the English lingua franca is a barrier for many developers [88, 99]. Furthermore, the incentives of OSS developers vary by geography: while developers in the USA show a relatively strong interest in “geek culture”, developers in India and China tend to be motivated primarily by career benefits [100]. Thus, “researchers studying open source should be mindful of geographic variation in what motivates participation and what forms participation may take, particularly outside of the code repository” [86].
Meanwhile companies primarily participate in OSS development for strategic reasons, such as recruiting developers [101, 102, 103], reducing costs [104, 105, 102], influencing OSS projects [106, 105], promoting open standards [104, 107, 108], and building a reputation as an OSS patron [90, 109, 32]. Commercial participation has mixed effects on the social structure of OSS communities. Typically, one company or a few companies emerge as dominant contributors in projects [28, 110]. The dominance of a company is negatively associated with the participation of volunteers, while it is positively associated with the productivity of contributors and the quality of issue reports [111, 74]. It is also common for companies, which may be market rivals, to collaborate in OSS ecosystems [112, 113, 110, 114, 115], which has turned many OSS communities “from networks of individuals into networks of companies” [101].
@@ 3.5 Data Analysis
To investigate development activity on the HF Hub (RQ1), we conducted a descriptive analysis of various types of activity in 348,181 model repositories, 65,761 dataset repositories, and 156,642 space repositories. Pearson correlation coefficients were calculated to assess the pairwise relationships between the activity variables. In addition, we employed the Mann-Whitney U test to compare the levels of activity across repositories with different licenses (Permissive, Restrictive, and No license). The Mann-Whitney U test is a non-parametric test that examines whether two independent samples come from the same distribution, which does not require the data to be normally distributed or to meet the assumption of homogeneity of variance [128]. Given the large sample sizes, the U values are expected to be large, and the salient test statistic is the p-value which indicates the statistical significance of observed differences. Due to capacity constraints in labelling licenses, we limited this analysis to repositories with licenses used in at least 100 repositories ($n=$339,502, 98% of all repositories). Subsequently, we analysed a snapshot of the social network structure of collaboration on the HF Hub (RQ2), using techniques defined in Table 1 in Appendix A. This analysis provides insights into collaboration patterns in model repositories at this point in time. Furthermore, we analysed collaboration patterns in the three AI sub-fields (NLP, CV, and MM) to enable comparisons. Lastly, we examined model adoption on the HF Hub (RQ3) by calculating the ranked degree of models in the bipartite model usage networks and ranked degree of models in the model co-usage networks to identify the most used models in spaces and their respective developers. These two complementary approaches quantified model popularity (i.e. which models are most frequently used in spaces) and model co-popularity (i.e. which models are most commonly used in conjunction with other models). We replicated this analysis for spaces with NLP, CV, and MM tags for comparative analysis of the three AI sub-fields.
@@ 4.1 Development Activity on the HF Hub
In this section, we present the findings of development activity in the repositories of 348,181 models, 65,761 datasets, and 156,642 spaces on the HF Hub. We present three key findings: right-skewed distributions across different types of activity (Section 4.1.1), strong correlations between development activities (Section 4.1.2), and a significant lack of licenses in model and dataset repositories (Section 4.1.3).
@@ 4.1.1 Right-Skewed Distributions in Development Activity
Activity per repository is extremely imbalanced, with right-skewed distributions of n_likes, n_discussions, n_commits, and n_downloads across model, dataset, and space repositories (see Figure 1). For example, while the maximum number of likes amongst models is over 9,000, the average model only receives 1.14 likes (see Tables 4-4). The majority of repositories get minimal engagement. For example, 91% of models and 88% of datasets have 0 likes; 84% of models, 91% of datasets, and 96% of spaces have 0 discussions; and 71% of models and 70% of datasets have 0 downloads. Meanwhile, most activity is concentrated in a small number of repositories. For example, $<1$% of models account for 80% of likes, 10% for 80% discussions, 30% for 80% commits, and $<1$% for 80% downloads. Upon increasing the threshold, 8% of models account for 99% likes, 15% for 99% discussions, and 1% for 99% downloads.
@@
Figure 1: Distributions of Development Activity in HF Hub Repositories
@@ 4.1.2 Correlation between Community Size and Engagement
4.1.3 Impact of Licenses on Collaboration
A significant proportion of model and dataset repositories lack licenses, which can create uncertainty and potential legal issues for users and developers. Specifying a license is not the norm: the majority of model repositories (65%) and datasets (72%) do not have a license. Amongst the licensed models, the most commonly used licenses are Apache v2.0 (37%), MIT (17%), OpenRAIL (14%), and CreativeML OpenRAIL-M (10%). The most used licenses for datasets are MIT (28%), Apache v2.0 (15%), OpenRAIL (9%), and licenses from the family of Creative Commons v4.0 (7%).
@@ 4.1.3 Impact of Licenses on Collaboration
Figure 3: Correlations of Activity in Model Repositories with Different Licenses
The choice of license matters: there is a moderate to strong correlation between the use of a license and level of activity in model repositories (see Figure 3). Furthermore, the Mann-Whitney U tests provide strong evidence of statistically significant differences between collaboration dynamics in model repositories with different types of licenses (all tests have $p<0.001$). Specifically, model repositories with permissive licenses consistently have the highest levels of activity compared to model repositories with no license and those with restrictive licenses (see Table 5). However, repositories with restrictive licenses also exhibit significantly higher activity than those with no license. This pattern holds across all activity metrics measured, suggesting that while permissive licenses foster the highest engagement, restrictive licenses also promote more collaboration compared to model repositories that do not have a license.
@@ 4.2.1 Collaboration in Model Repositories on the HF Hub
Collaboration is characterised by high reciprocity values, ranging from 0.81 to 1.00 across all k-core levels (see Table 9), indicating the prevalence of mutual relationships amongst developers. The low assortativity values, ranging from -0.49 to 0.08, suggest that developers collaborate regardless of their centrality in the network, implying that other factors, such as shared interests, skills, or project roles, may be more significant in driving collaboration than their network centrality. Furthermore, the relatively low average rich club coefficients, ranging from 0.04 to 0.41, indicate that highly central developers do not primarily collaborate with each other and a lack of elitism amongst power developers.
@@ 5.1.1 Contributions to Academic Literature
Impact of license on collaboration: The Mann-Whitney U tests show that license choice significantly impacts the level of activity and engagement in repositories, with permissive licenses exhibiting the highest activity levels, followed by repositories with restrictive licenses, and finally ones with no license. Furthermore, the Pearson correlations indicate that the use of a license (permissive or restrictive) is associated with stronger correlations between various types activity compared to repositories without licenses. These findings highlight the important role of licencing decisions in influencing the collaborative and community dynamics in open model development and open source AI projects.
@@ 5.1.3 Recommendations for Future Research
1. Implications of concentrations in the HF Hub developer community: We confirm prior observations that the models of a handful of companies are dominant amongst the HF Hub developer community [21]. We encourage future research to investigate what these concentrations mean in practice, such as the potential benefits that these companies accrue from their open model ecosystems, including increased visibility, crowdsourced contributions (e.g., via commits and discussions), and access to diverse fine-tuned versions shared by other developers on the HF Hub. Furthermore, there is a concern that dominant companies benefit from developers being locked-in to their ecosystems, potentially limiting competition and entrenching their dominance. Future research could investigate the factors contributing to such concentrations, such as the reputation of the companies developing the models, their access to resources and support, or the perceived performance and versatility of their models, as well as the implications of these concentrations for the broader AI community, including the impact on research, innovation, and the distribution of benefits and resources.
@@
2. Incentives and modes of participation: Future research could investigate the incentives of individual developers and companies. A number of companies have released open models on the HF Hub, such as Meta’s LlaMA models [129], Mistral AI’s Mixtral models [130], and OpenAI’s Whisper models [131]. Often these releases are presented as acts of “AI democratisation” [132]. Future research could critically examine the commercial incentives behind these releases. In addition, future research could examine commercial approaches to model governance and maintenance—for example, if and how companies welcome or engage with community contributions—and if and how companies collaborate with each other on open model development, as they do in OSS development [110, 112, 113, 114, 115].
3. Collaboration dynamics in active repository communities: We know that model maintenance focuses on model performance improvements [22]; and in the minority of repositories that have active communities, most developers contribute to discussions rather than commits (see Tables 4-4). Going further, we encourage researchers to examine collaboration dynamics in repositories with active communities from multiple angles. Given the sizeable differences in n_committers and n_disc_starters, future research could investigate the division of roles between discussion and code contributors, typical topics of discussion (e.g., model performance, new ideas, etc.), how discussions inform model maintenance if at all, and the journeys of developers from discussion contributors to committers, amongst others. In addition, future research could examine the governance approaches (e.g., contribution policies) that repository owners use to encourage collaboration. Future analyses could also take into account temporal dynamics, providing insights into evolving patterns, social structures, and trends of open model developer communities on the HF Hub.
@@
5. Collaboration beyond the HF Hub: While this analysis provides insights into the developer community that shares and fine-tunes models on the HF Hub, we have a limited understanding of the development of the various components involved in the development of models [4], which largely takes place in proprietary settings or on other platforms like GitHub [14]. We encourage future research to examine how the HF Hub is used in the wider ecosystem of platforms and offline venues for the collaborative development of open models and datasets. This research direction would enable comparisons of the collaboration patterns amongst model developers and model fine-tuners. In addition, researchers could undertake a multi-sited analysis, examining collaboration on the same project across platforms.